Suppose marginal utility of good X is 20 while its price is Rs. 4 per unit and marginal utility of



Y good is 50 while its price is Rs.5 per unit .The individual to whom this information applies is



spending 20 on each good .Is he maximizing his satisfaction.?

1
Expert's answer
2021-12-03T08:36:33-0500

Solution:

 

Profit maximization: "\\frac{MU_{x} }{MU_{y} } = \u200b\n\\frac{Px}{Py }"

 

"\frac{20}{50 } = \frac{4}{5 }"

0.4 = 0.8

 

Budget constraint: I = PxX + PyY

 

20 = 4X + 5Y

 

The individual is not maximizing his satisfaction since there is no equality: